OpenStudy (anonymous):

WHAT'S THE DIFFERENCE BETWEEN Y= AND f(X)=..? I AM WORKING ON A MATH PACKET AND CAME ACROSS THIS; Y= -x^2 + 2x + 3 AND f(x)= -x^2 + 2x +3, I HAVE TO GRAPH BOTH OF THESE, MY QUESTION IS, WHAT IS THE DIFFERENCE BETWEEN BOTH OF THESE GRAPHS?

OpenStudy (anonymous):

._. we don't like ape we consider it as yelling

OpenStudy (jdoe0001):

ape? actually, I don't mind apes, they're our cousins primates :)

OpenStudy (anonymous):

y and f(x) are the same thing f(x) means function of x

OpenStudy (jdoe0001):

if you meant this "what's the difference between y= and f(x)=..? i am working on a math packet and came across this; y= -x^2 + 2x + 3 and f(x)= -x^2 + 2x +3, i have to graph both of these, my question is, what is the difference between both of these graphs? " in lowercase, mind the caps btw

OpenStudy (jdoe0001):

f(x) is a way to say... " a function whose value depends on the variable x in it" a short way to variablize it is just writing it as the variable "y" same thing